# Blind Spot Navigation in LLM Reasoning with Thought Space Explorer

### v2

@@ Abstract
Recent advances in large language models (LLMs) have demonstrated their potential in handling complex reasoning tasks, which are usually achieved by constructing a thought chain to guide the model to solve the problem with multi-step thinking. However, existing methods often remain confined to previously explored solution spaces and thus overlook the critical blind spot within LLMs’ cognitive range. To address these issues, we design the Thought Space Explorer (TSE), a novel framework to expand and optimize thought structures to guide LLMs to explore their blind spots of thinking. By generating new reasoning steps and branches based on the original thought structure with various designed strategies, TSE broadens the thought space and alleviates the impact of blind spots for LLM reasoning. Experimental results on multiple levels of reasoning tasks demonstrate the efficacy of TSE. We also conduct extensive analysis to understand how structured and expansive thought can contribute to unleashing the potential of LLM reasoning capabilities.
Blind Spot Navigation in LLM Reasoning with Thought Space Explorer
Jinghan Zhang1, Fengran Mo2, Xiting Wang3, Kunpeng Liu1thanks: Corresponding Author. 1Portland State University, 2University of Montreal, 3Renmin University of China {jinghanz,kunpeng}@pdx.edu fengran.mo@umontreal.ca xitingwang@ruc.edu.cn
1 Introduction
@@
Recent advances in large language models (LLMs) have shown great potential in solving complex tasks with reasoning capabilities Huang and Chang (2022); Achiam et al. (2023); Patterson et al. (2022); Mao et al. (2023) by guiding the LLMs to reasoning logically solve the complex task step-by-step. A common practice is to design the Chain-of-Thought (CoT) Kojima et al. (2022) to boost reasoning capabilities by evolving the thinking from single outputs to intermediate steps reasoning.
@@
Existing studies Wang et al. (2022); Yao et al. (2024); Zhang et al. (2024d); Besta et al. (2024) attempt to develop various thought structures with multiple chains or branches of thought on top of CoT to arouse the LLM reasoning ability. Compared with direct output and CoT, the core advantage of thought structures enables models to explore the solution space of a task from local to global Hao et al. (2023). For example, in Figure 1, thought structures may initiate exploration from two distinct points “specialty” and “industry”. Such exploration allows LLMs to generate diverse paths to solutions and thus enhances reasoning capacity. Moreover, the diverse structures can enable models to perform forward and backward evaluations within the explored thought space to seek the optimal solution, i.e., a more effective reasoning thought.
Figure 1: Thought structure optimization through TSE. On the left side, we showcase traditional thought structures and optimization methods, where the LLM’s generation may limited by its thought pattern. On the right side, we show how TSE method expands thought structure through a three-step generation of branches. Our method guides the LLM to explore the blind spots between previous thought paths.
@@
A series of studies are conducted to optimize thought structures with various aspects, including generating parallel thought Wang et al. (2022), constructing tree structure reasoning topology on top of CoT Yao et al. (2024), fine-tuning the LLMs with direct preference optimization (DPO) to align thought steps of CoT with task performance Zhang et al. (2024c), etc. The key idea of this work is to compare multiple responses or extending existing chains (e.g. “Coffee over Drone” or “Coffee industry $\rightarrow$ Coffee bottle industry”) to obtain a better thought chain. However, these approaches lack exploration of unknown solution spaces, which might lead to consistent oversight of LLMs’ cognitive blind spots Liu et al. (2024a); Zhang et al. (2024a); Sprague et al. (2024a); Mo et al. (2024). Merely generating more chains does not enable LLMs to conceive of content previously unthought of. This is because the over-generated chains tend to repeat previous thought patterns and would lead to two main issues: (1) The absence of feasible solutions, when feasible solutions are in blind spot space, repeatedly filtering or extending existing paths may converge to a local optimum (e.g., exploring only from a coffee perspective). (2) Insufficient diversity. Especially for open-ended questions, existing methods have limited impact on exploring the thought space, and excessive extension or filtering might even reduce the diversity of responses (e.g., discarding feasible solutions or creating redundancy through repetitive thinking).
@@
To address these issues, we propose the Thought Space Explorer (TSE), a novel framework designed to expand and optimize thought structures. The TSE starts from already explored thought paths and guides the model to explore hidden solution spaces because the existing thought structures often already contain feasible solutions or crucial information pointing towards such solutions. To enhance efficiency and precision, further exploration of the model starts from thought nodes within explored solutions, which ensures that the reasoning process is not a blind exploration but a deeper inquiry based on verified insights.
@@
Concretely, to identify key points of information from existing thoughts, we first quantify the contribution of each thought node to the conclusion during the model’s reasoning process to select key thought nodes (e.g., in Chain 2, the details about “drones and delivery” serve as key information leading toward “logistics industry”). Considering the visibility of parameters in LLMs, we select key nodes from two perspectives, relative gradients and semantic relationships. Further, based on these key nodes, the model generates new thought nodes and proceeds with deeper reasoning in new directions from “original node” to “new nodes” and facilitates the exploration of the solution space by the thought structure. Finally, depending on whether the reasoning tasks require a singular or comprehensive conclusion, we proceed with collaborative reasoning across the entire thought structure to generate the output. We evaluate the effectiveness of TSE on three reasoning tasks and the results show that TSE significantly improves the performance of thought structures compared with existing methods. We further analyze the effectiveness of each component to comprehensively understand their functionality in each stage.
@@
Our contributions are summarized as follows: (1) We propose TSE reasoning framework to expand thought structures for exploring solution spaces to alleviate the impact of blind spots for LLM reasoning. (2) We investigate various strategies to prioritize and refine the thought structure by identifying the importance of the node in the thought structure. Different strategies are designed to adapt to various settings. (3) Experimental results on three specific reasoning tasks indicate the effectiveness of our TSE compared with the existing reasoning methods without exploring thought structure.
2 Related Work
2.1 LLM Reasoning Structures
@@
The most straightforward method to address reasoning tasks is to generate a conclusion through one-step thinking. However, the LLM might overlook essential intermediate steps and generate incoherent logic or incorrect conclusions Chu et al. (2023). The advent of CoT Wei et al. (2022); Wang and Zhou (2024) optimizes the reasoning step by connecting distinct thoughts into a coherent sequence Li et al. (2024). Although CoT can improve transparency and coherence, its singular structure limits its capability to handle more complex logical relationships Jin and Lu (2024); Sprague et al. (2024b). To this end, some studies develop structured reasoning methods, such as self-consistent CoT and Tree-of Thought. Wang et al. (2022); Zhang et al. (2024d); Yao et al. (2024); Liu et al. (2024b); Mo and Xin (2024); Zhang et al. (2024b, a). These sophisticated thought structures enhance the consistency and systematic nature of reasoning by expanding the ability of the model to manage diverse logical relationships Xia et al. (2024); Stechly et al. (2024). Thought structures offer distinct advantages by maintaining coherence and depth while increasing the diversity and flexibility of reasoning pathsLiang et al. (2024). However, the reasoning chains within these structures are highly repetitive and thus reduce generation and selection efficiency.
2.2 Thought Structure Optimizations
@@
To further enhance the capabilities of thought structures, recent researches focus on two main optimization strategies. The first is selecting optimal paths within the structures Feng et al. (2023); Long (2023); Hao et al. (2023); Shinn et al. (2023); Jung et al. (2022). By choosing optimal paths, the model filters the irrelevant and low-quality branches and globally searches for correct or optimal solutions, thus enhancing reasoning efficiency and accuracy. The second is expanding reasoning depth and breadth Zhu et al. (2022); Besta et al. (2024); Gao et al. (2024); Zhang et al. (2024b, c); Hou et al. (2024). By deepening and widening the thought structure, the model can explore a broader array of possibilities and perspectives, thus improving its understanding and capability to solve complex issues. However, these methods might be limited to previously explored spaces and directions, thus failing to adequately investigate the blind spots within the thought space of the model. Different from them, we aim to expand the depth and breadth of thought structures by actively exploring the cognitive blind spots of the model in thought space.
3 Methodology
@@
To expand and optimize thought structures for reasoning tasks, we introduce TSE, a self-expanding and exploring method to enable LLMs to proactively address deficiencies in reasoning processes and explore new reasoning directions with limited generation. We implement the TSE method through several steps: (1) Key Node Selection, in which we identify the most influential nodes and generate new nodes based on the crucial information they contain; (2) Connection and Expansion, in which we systematically connect selected key nodes and expands them into new branches to explore new reasoning directions; and (3) Collaborative Reasoning, in which we address deficiencies in the model’s ability to synthesize and integrate diverse reasoning paths in different directions.
3.1 Problem Formulation
@@
Given a specific reasoning task $\mathcal{Q}$, a large language model (LLM) $\mathcal{L}$ is expected to generate a thought structure $\mathcal{S}$ composed of directed links connecting sentence-level thought nodes. The set of all possible thought nodes, e.g., sentences or reasoning steps, is denoted as $\mathcal{T}$, where $T_{ij}$ denotes the $j$-th thought point in the $i$-th chain. Then the thought point $\mathcal{T}$ and the connections between two nodes compromise the set of vertices $\mathbf{V}$ and the directed edges $\mathbf{E}$ in the thought structure $\mathcal{S}$. Consequently, a thought chain $C_{i}$ is an ordered sequence of thought nodes:
$\mathbf{V}=\bigcup_{i=1}^{N}\bigcup_{j=1}^{K_{i}}\{T_{ij}\},K_{i}=|C_{i}|,$ (1)
$\mathbf{E}=\bigcup_{i=1}^{N}\bigcup_{j=1}^{K_{i}-1}\{(T_{ij},T_{i(j+1)})\},$ (2)
@@
Figure 2: Solution space exploration by TSE. By generating new branches of solutions, the explored space of solution expands.
For a specific task $\mathcal{Q}$, the complete reasoning solution space $\mathcal{P}$ encompasses all possible valid reasoning paths (thought chains). As shown in Figure 2, the space that has been explored by the generated thought structure $\mathcal{S}$ and the unexplored implicit thought space are denoted as $\mathcal{P_{S}}$ and $\mathcal{P_{U}}$, where $\mathcal{P_{S}}\cup\mathcal{P_{U}}=\mathcal{P}$. Our objective is to generate new branches $C^{\prime}$ and expand $\mathcal{S}$ to cover as many as possible $\mathcal{P_{U}}$ to increase the likelihood of finding viable solutions, and the optimization goal is defined as Eq. 4, where $J$ is the reasoning performance metric and $\mathcal{S}^{\prime}$ is the expanded thought structure.
$\max_{\mathcal{S}^{\prime}}J(\mathcal{S}^{\prime},\mathcal{Q})$ (4)
3.2 Key Node Selection
The key nodes refer to those significantly impacting the solution path in the solution space $\mathcal{P}$, i.e., containing the known crucial information required to solve the problem. We aim to select key nodes within $\mathcal{S}$, a structure composed of multiple parallel thought chains with the same length. Then, the thought chains $\mathcal{S}$ can be expanded based on the selected key nodes from two aspects: i) exploring the most promising areas for solution paths, and ii) reducing the error propagation by extra analysis and checking on the key nodes, which are often the source of the possible errors. To support the different availability of the backbone models, we propose two methods, gradient-based and self-prompting selection to identify the key nodes for those with accessible gradients and black-box ones, respectively.
3.2.1 Gradient-based Selection
The Gradient-based selection is applicable when the internal structure and the gradient information of model $\mathcal{L}$ is accessible. The representation of each thought node $T_{ij}$ in the generated thought chains is obtained by the model $\mathcal{L}$ as:
$\mathcal{L}:T_{ij}\rightarrow\mathbf{v}_{ij}\in\mathbb{R}^{d}$ (5)
Then, we aim to analyze the gradient importance of nodes $\mathbf{T}_{ij}$ relative to the conclusion node $T_{iK_{i}}$ within a chain. The representation of the conclusion node $v_{iK_{i}}$ is mapped to the output space as the model’s prediction at the end of the chain:
@@ 3.2.1 Gradient-based Selection
where the function $f$ represents a mapping from the representation space of the conclusion node $\mathbf{v}_{iK_{i}}$ to the output space, and the prediction $\hat{y}_{i}$ is usually a textual conclusion or decision to task $\mathcal{Q}$.
The self-information loss $L_{i}$ is a common practice to evaluate the model’s confidence in its predictions Wang and Feng (2021), where higher confidence corresponds to lower loss values. Thus, we calculate the partial derivative of the loss $\mathbf{g}_{ij}$ with respect to each node’s representation $\mathbf{v}_{ij}$ and the Euclidean norm of its gradient $G_{ij}$ to measure the importance of the nodes. Then, a normalization is applied to determine the relative importance $I_{ij}$ of each node for a consistent and comparative analysis of node importance across different chains within the whole thought structure:
$L_{i}=-\log P(\hat{y}_{i}\mid\mathbf{v}_{iK_{i}}),$ (7)
$\mathbf{g}_{ij}=\frac{\partial L_{i}}{\partial\mathbf{v}_{ij}},\quad G_{ij}=\|\mathbf{g}_{ij}\|_{2},$ (8)
$I_{ij}=\frac{G_{ij}}{\sum_{k=1}^{K_{i}}G_{ik}}$ (9)
In this way, we can identify the key nodes that have the highest impact on the model’s predictions. Finally, a set of key nodes $\mathcal{T}_{key}=\left\{T_{i_{key}}\right\}_{i=1}^{N}$ is obtained with the highest relative gradient.
3.2.2 Self-prompting Selection
When the gradient is not accessible, e.g., a black-box model, the self-prompt selection is an alternative for key node identification. Under this setting, we leverage LLM’s natual language understanding capabilities to analyze and prioritize key nodes in the thought chains. Although the inner workings of the model are opaque, we infer critical areas within the network’s structure by constructing specific prompts based on semantic and logical relationships. Specifically, model $\mathcal{L}$ ranks the importance of nodes within chain $C_{i}$ for the key node as:
@@ 3.3 Connection and Expansion
After obtaining selected key nodes, the next step is to adopt them as conditional information to generate new thought nodes. Then we generate the new nodes based on the key nodes by integrating them into the thought structure and expanding it to form new branches. For each new node, we select two key nodes from $\mathcal{T}_{key}$, denoted as $T_{i_{key}}$ and $T_{l_{key}}$. With each pair of selected key nodes, the model generates a new thought node $T^{1}_{il}$ as
$T^{1}_{il}=\mathcal{L}(T_{i_{key}},T_{l_{key}}),i,l\in[1,N],i\neq l$ (11)
@@
For the new node, to decide which key node to follow, we need to select $T_{i_{key}}$ or $T_{l_{key}}$ to serve as the connection point $T^{1}_{il}$ for extending the new thought branch, which is supposed to maintain logical coherence while exploring previously unknown regions of the thought space. Therefore, we select the conditional nodes with higher semantic relevance with new nodes and contribute more to reasoning as connection nodes. Building on the previous stage, we propose two connection node selection methods: the relative gradient method and the semantic relevance method.
3.3.1 Relative Gradient Selection
@@
The relative gradient selection is applied to take the node between $T_{i_{key}}$ and $T_{l_{key}}$ with the larger relative gradient as the connection point by comparing the importance indices $I_{i_{key}}$ and $I_{l_{key}}$:
$C^{\prime}=\langle\arg\max_{T_{key}\in\{T_{i_{key}},T_{l_{key}}\}}(I_{key}),T^{1}_{il}\rangle$ (12)
@@
The $C^{\prime}$ denotes the new branch initiated from the key node with the higher importance index. Starting from $T^{1}_{il}$, the model $\mathcal{L}$ continues reasoning guided by the information from the new nodes, and since the new nodes introduce fresh perspectives of existing information, the new branch is likely to go and explore previously unconsidered directions with step-by-step reasoning. The process of extending the branch $C^{\prime}$ continues until a specified depth is reached. The language model $\mathcal{L}$ recursively generates subsequent nodes, using the strongest connection node as the context for generating new nodes:
$C^{\prime}\rightarrow\mathcal{L}(C^{\prime})=\langle C^{\prime},T^{2}_{il},\ldots,T^{K}_{il}\rangle,$ (13)
$K=\begin{cases}K_{i}&\text{if }I_{i_{key}},>I_{l_{key}}\\ K_{l}&\text{otherwise}.\end{cases}$ (14)
3.3.2 Semantic Relevance Selection
@@
Semantic relevance selection is applicable after both key node selection methods. Here, the connection node is selected by the model $\mathcal{L}$ based on semantic relationships between nodes. Then the new branch $C^{\prime}$ continues to extend recursively until a specified depth is reached, with $\mathcal{L}$ generating subsequent nodes based on the strongest semantic relationship:
@@
$C^{\prime}\rightarrow\mathcal{L}(C^{\prime})=\left\langle T_{\text{selected}},T_{il}^{1},T_{il}^{2},\ldots,T_{il}^{K}\right\rangle$ (15)
@@
where $T_{selected}$ is the key node chosen by $\mathcal{L}$, and $K$ denotes the ending depth of the thought chain.
3.4 Collaborative Reasoning
@@
Give the task $\mathcal{Q}$ and its unseen complete solution space $\mathcal{P}$, the model needs to generate new thought branches based on the original thought structure $\mathcal{S}$. During this process, each new branch $C^{\prime}$ expands the explored thought space $\mathcal{P_{S}}\in\mathcal{P}$ by mining potential solutions based on the established structure:

### v3

# Blind Spot Navigation in Large Language Model Reasoning with Thought Space Explorer

@@ Abstract
Recent advances in large language models (LLMs) have demonstrated their potential in handling complex reasoning tasks, which are usually achieved by constructing a thought chain to guide the model in solving the problem with multi-step thinking. However, existing methods often remain confined to previously explored solution spaces and thus overlook the critical blind spot within LLMs’ cognitive range. To address these issues, we introduce the “Thought Space Explorer” (TSE), a novel framework to expand and optimize thought structures to guide LLMs to explore their blind spots of thinking. By generating new reasoning steps and branches based on the original thought structure with various designed strategies, TSE broadens the thought exploration view and alleviates the impact of blind spots for LLM reasoning. Experimental results on multiple levels of reasoning tasks demonstrate the efficacy of TSE by surpassing various baseline methods. We also conduct extensive analysis to understand how structured and expansive thought can contribute to unleashing the potential of LLM reasoning capabilities.
Introduction
@@
Recent advances in large language models (LLMs) have shown great potential in solving complex tasks with reasoning capabilities (Huang and Chang 2022; Patterson et al. 2022; Achiam et al. 2023; Mao et al. 2023; Dutta et al. 2025) by guiding the LLMs to reasoning logically solve the complex task step-by-step. A common practice is to design the Chain-of-Thought (CoT) (Kojima et al. 2022) to boost reasoning capabilities by evolving the thinking from single outputs to intermediate steps reasoning.
@@
Existing studies (Wang et al. 2022; Yao et al. 2024; Zhang, Yuan, and Yao 2024; Besta et al. 2024; Pandita et al. 2025) attempt to develop various thought structures with multiple chains or branches of thought on top of CoT to arouse the LLM reasoning ability. Compared with direct output and CoT, the core advantage of thought structures enables models to explore the solution space of a task from local to global (Hao et al. 2023). For example, in Figure 1, thought structures may initiate exploration from two distinct points “specialty” and “industry”. Such exploration allows LLMs to generate diverse paths to solutions and thus enhances reasoning capacity. Moreover, the diverse structures can enable models to perform forward and backward evaluations within the explored thought space to seek the optimal solution, i.e., a more effective reasoning thought.
A series of studies are conducted to optimize thought structures with various aspects, including generating parallel thought (Wang et al. 2022), constructing tree structure reasoning topology on top of CoT (Yao et al. 2024), fine-tuning the LLMs with direct preference optimization (DPO) to align thought steps of CoT with task performance (Zhang et al. 2024c), etc. The key idea of this work (see Figure 1) is to compare multiple responses or extending existing chains (e.g. “Coffee over Drone” or “Coffee industry $\rightarrow$ Coffee bottle industry”) to obtain a better thought chain.
@@
However, these approaches lack exploration of unknown solution spaces, which might lead to consistent oversight of LLMs’ cognitive blind spots (Zhang et al. 2024a; Sprague et al. 2025; Mo et al. 2024; Liu et al. 2025; Weerasooriya et al. 2023). Merely generating more chains does not enable LLMs to conceive of content previously unthought of. As described in Figure 1, over-generated chains tend to repeat prior thought patterns, leading to two main issues: (1) the absence of feasible solutions—when such solutions lie in blind spot regions, repeatedly filtering or extending existing paths may converge to a local optimum (e.g., exploring only from a coffee perspective); and (2) insufficient diversity—especially for open-ended questions, where existing methods have limited impact on exploring the thought space, and excessive extension or filtering might even reduce the diversity of responses (e.g., discarding feasible solutions or creating redundancy through repetitive thinking).
@@
To address these issues, we propose the Thought Space Explorer (TSE), a novel framework designed to expand and optimize thought structures. The TSE starts from thought paths already explored and guides the model to explore hidden solution spaces because the existing thought structures often already contain feasible solutions or crucial information pointing towards such solutions. To enhance efficiency and precision, further exploration of the model starts from thought nodes within explored solutions, which ensures that the reasoning process is not a blind exploration but a deeper inquiry based on verified insights.
Figure 1: Thought structure optimization through TSE. On the left side, we showcase traditional thought structures and optimization methods, where the LLMs’ generation may limited by its thought pattern. On the right side, we show how TSE expands thought structure through a three-step generation of branches. TSE guides LLMs to explore the blind spots between previous thought paths.
@@
To identify key points of information from existing thoughts, as shown in Figure 1, we first quantify each thought node’s contribution to the conclusion during the model’s reasoning process to select key nodes (e.g., in Chain 2, the details about “drones and delivery” serve as key information leading toward “logistics industry”). Considering the visibility of parameters in LLMs, we select key nodes from two perspectives: relative gradients and semantic relationships. Based on these key nodes, the model then generates new thought nodes and proceeds with deeper reasoning in new directions from “original nodes” to “new nodes”, facilitating exploration of the solution space through the thought structure. Finally, depending on whether the reasoning task requires a singular or comprehensive conclusion, we perform collaborative reasoning across the entire thought structure to generate the output.
We evaluate the effectiveness of TSE on three reasoning tasks and the results show that TSE significantly improves the performance of thought structures compared with existing methods. We further analyze the effectiveness of each component to comprehensively understand their functionality in each stage.
@@
Our contributions include: (1) We propose TSE reasoning framework to expand thought structures for exploring solution spaces to alleviate the impact of blind spots for LLM reasoning. (2) We investigate various strategies to prioritize and refine the thought structure by identifying the importance of the node in the thought structure. Different strategies are designed to adapt to various settings. (3) Experimental results on three specific reasoning tasks indicate the effectiveness of our TSE compared with the existing reasoning methods without exploring thought structure.
Related Work
LLM Reasoning Structures
@@
The most straightforward method to address reasoning tasks is to generate a conclusion through one-step thinking. However, the LLM might overlook essential intermediate steps and generate incoherent logic or incorrect conclusions (Chu et al. 2023). The advent of CoT (Wei et al. 2022; Wang and Zhou 2024) optimizes the reasoning step by connecting distinct thoughts into a coherent sequence (Li et al. 2024). Although CoT can improve transparency and coherence, its singular structure limits its capability to handle more complex logical relationships (Jin and Lu 2024; Sprague et al. 2025). To this end, some studies develop structured reasoning methods, such as self-consistent CoT and Tree-of Thought. (Wang et al. 2022; Zhang, Yuan, and Yao 2024; Yao et al. 2024; Liu et al. 2024; Mo and Xin 2024; Zhang et al. 2024b, a). These sophisticated thought structures enhance the consistency and systematic nature of reasoning by expanding the ability of the model to manage diverse logical relationships (Xia et al. 2024; Stechly, Valmeekam, and Kambhampati 2024). Thought structures offer distinct advantages by maintaining coherence and depth while increasing the diversity and flexibility of reasoning paths (Liang et al. 2024). However, the reasoning chains within these structures are highly repetitive and thus reduce generation and selection efficiency.
Thought Structure Optimizations
@@
To further enhance the capabilities of thought structures, recent researches focus on two main optimization strategies. The first is selecting optimal paths within the structures (Feng et al. 2023; Long 2023; Hao et al. 2023; Shinn, Labash, and Gopinath 2023; Jung et al. 2022). By choosing optimal paths, the model filters the irrelevant and low-quality branches and globally searches for correct or optimal solutions, thus enhancing reasoning efficiency and accuracy. The second is expanding reasoning depth and breadth (Zhu et al. 2022; Besta et al. 2024; Gao et al. 2024; Zhang et al. 2024b, c; Hou et al. 2024; Zhang et al. 2025). By deepening and widening the thought structure, the model can explore a broader array of possibilities and perspectives, thus improving its understanding and capability to solve complex issues (Chen et al. 2025; Pan et al. 2025; Wang et al. 2025). However, these methods might be limited to previously explored spaces and directions, thus failing to adequately investigate the blind spots within the thought space of the model. Different from them, we aim to expand the depth and breadth of thought structures by actively exploring the cognitive blind spots of the model in thought space.
Methodology
@@
To expand and optimize thought structures for effective exploration of reasoning spaces, we introduce TSE, a self-expansion and exploration method that allows language models to proactively address deficiencies in reasoning processes and explore new reasoning directions with limited steps of generation. In a step-by-step reasoning scenario, we implement the TSE method through several steps: (1) Key Node Selection and Generation, in which we select the nodes that are most influential to exploration directions based on the crucial information contained previously, then a new node is generated to integrate the insights of the key node for new exploration directions; (2) Connection and Expansion, in which we connect and expand the exploration to potential new directions within the semantic context from the newly-generated node; and (3) Collaborative Reasoning, in which we address deficiencies in the model’s ability to synthesize and integrate diverse reasoning paths in different directions.
Problem Formulation
@@
Given a specific reasoning task $\mathcal{Q}$, we apply a large language model (LLM) $\mathcal{L}$ to a structured reasoning process as a thought structure $\mathcal{S}$. This structure consists of multiple reasoning sentences as thought nodes that are connected sequentially. We define the set of all thought nodes as $\mathcal{T}$, where each node $\mathcal{T}_{ij}$ represents the $j$-th reasoning step in the $i$-th thought chain. The thought structure $\mathcal{S}$ can be viewed as a directed graph consisting of vertices (thought nodes) $\mathbf{V}$ and edges (connections between consecutive nodes) $\mathbf{E}$, and a thought chain $C_{i}$ as an ordered sequence of thought nodes as
$\mathbf{V}$ $=\bigcup_{i=1}^{N}\bigcup_{j=1}^{K_{i}}\{T_{ij}\},\quad K_{i}=|C_{i}|,$ (1)
$\mathbf{E}$ $=\bigcup_{i=1}^{N}\bigcup_{j=1}^{K_{i}-1}\{(T_{ij},T_{i(j+1)})\}$
where $N$ and $K$ is the number of the thought chain $C$ and the number of the node $\mathbf{V}$ contained in a thought chain $C$, respectively.
@@
Figure 2: Illustration of solution space exploration via our TSE method. By generating new branches of solutions, the explored space of solutions expands.
For a specific task $\mathcal{Q}$, the complete reasoning solution space $\mathcal{P}$ encompasses all possible valid reasoning paths $C$ (thought chains). As shown in Figure 2, the space that has been explored by the generated thought structure $\mathcal{S}$ and the unexplored implicit thought space are denoted as $\mathcal{P_{S}}$ and $\mathcal{P_{U}}$, where $\mathcal{P_{S}}\cup\mathcal{P_{U}}=\mathcal{P}$.
Our goal is to actively expand the thought structure $\mathcal{S}$ by generating new reasoning branches $C^{\prime}$ to explore previously untouched areas of the solution space ($\mathcal{P_{U}}$). In this way, we maximize the chance of discovering effective and innovative solutions. The optimization goal is defined as $\max_{\mathcal{S}^{\prime}}J(\mathcal{S}^{\prime},\mathcal{Q})$, where $J$ is the reasoning performance metric and $\mathcal{S}^{\prime}$ is the expanded thought structure.
Key Node Selection and Generation
First, we select the most impactful nodes from the existing thought structure $\mathcal{S}$ for expansion. Specifically, we expect the nodes to contain crucial information necessary for solving the given problem by achieving the following aspects: (i) effective exploration of promising regions of the solution space by starting from these key nodes, thereby increasing the likelihood of discovering viable solutions; and (ii) minimization of error propagation by conducting additional analysis and verification on these critical nodes, which often represent potential sources of mistakes. To support the different availability of the backbone models, we propose two methods, gradient-based and self-prompting selection to identify the key nodes for those with accessible gradients and black-box ones, respectively.
Gradient-based Selection
The gradient-based selection is applicable when the internal structure and the gradient information of the model $\mathcal{L}$ are accessible. The representation of each thought node $\mathcal{T}_{ij}$ in the generated thought chains is obtained via the model $\mathcal{L}$ as:
$\mathcal{L}:\mathcal{T}_{ij}\rightarrow\mathbf{v}_{ij}\in\mathbb{R}^{d}$ (3)
Then, we aim to analyze the gradient importance of nodes $\mathcal{T}_{ij}$ relative to the conclusion node $\mathcal{T}_{iK_{i}}$ within a chain. The representation of the conclusion node $\mathbf{v}_{iK_{i}}$ is mapped to the output space as the model’s prediction at the end of the chain:
@@
where the function $f(\cdot)$ represents a mapping from the representation space of the conclusion node $\mathbf{v}_{iK_{i}}$ to the output space, and the prediction $\hat{y}_{i}$ is usually a textual conclusion or decision in terms of task $\mathcal{Q}$.
The self-information loss $L_{i}$ is a common practice to evaluate the model’s confidence in its predictions (Wang and Feng 2021), where higher confidence corresponds to lower loss values. Thus, we calculate the partial derivative of the loss $\mathbf{g}_{ij}$ with respect to each node’s representation $\mathbf{v}_{ij}$ and the Euclidean norm of its gradient $G_{ij}$ to measure the importance of the nodes. Then, we apply a normalization to determine the relative importance $I_{ij}$ of each node for a consistent and comparative analysis of node importance across different chains within the thought structure $\mathcal{S}$:
$L_{i}=-\log P(\hat{y}_{i}\mid\mathbf{v}_{iK_{i}}),\quad\mathbf{g}_{ij}=\frac{\partial L_{i}}{\partial\mathbf{v}_{ij}}$ (5)
$\quad G_{ij}=\|\mathbf{g}_{ij}\|_{2},\quad I_{ij}=\frac{G_{ij}}{\sum_{k=1}^{K_{i}}G_{ik}}$ (6)
In this way, we can identify the key nodes that have the highest impact on the model’s predictions. Now we have obtained a set of key nodes $\mathcal{T}_{key}=\left\{T_{i_{key}}\right\}_{i=1}^{N}$ with the highest relative gradient.
Self-prompting Selection
When the gradient is not accessible, e.g., a black-box model, the self-prompt selection is an alternative for key node identification. Under this setting, we leverage LLM’s natual language understanding capabilities to analyze and prioritize key nodes in the thought chains. Although the inner mechanisms of the model are opaque, we infer critical areas within the network’s structure by constructing specific prompts based on semantic and logical relationships. Specifically, model $\mathcal{L}$ ranks the importance of nodes within chain $C_{i}$ for the key node as:
@@
With the selected key nodes, the next step is to adopt them as conditional information to generate new thought nodes. Then we generate the new nodes based on the key nodes by integrating them into the thought structure and expanding them to form new branches. For each new node, we select two key nodes from $\mathcal{T}_{key}$, denoted as $\mathcal{T}_{i_{key}}$ and $\mathcal{T}_{l_{key}}$. With each pair of selected key nodes, the model generates a new thought node $\mathcal{T}^{1}_{il}$ as:
$\mathcal{T}^{1}_{il}=\mathcal{L}(\mathcal{T}_{i_{key}},\mathcal{T}_{l_{key}}),i,l\in[1,N],i\neq l$ (8)
Connection and Expansion
@@
In terms of a new node, to decide which key node it should follow, we need to select $\mathcal{T}_{i_{key}}$ or $\mathcal{T}_{l_{key}}$ to serve as the connection point $\mathcal{T}^{1}_{il}$ for extending the new thought branch, which is supposed to maintain logical coherence while exploring previously unknown regions of the thought space. Therefore, we select the conditional nodes with higher semantic relevance with new nodes and contribute more to reasoning as connection nodes. Building on the previous stage, we propose two connection node selection methods, including the relative gradient method and the semantic relevance method.
Relative Gradient Selection
@@
The relative gradient selection is applied to take the node between $\mathcal{T}_{i_{key}}$ and $\mathcal{T}_{l_{key}}$ with the larger relative gradient as the connection point by comparing the importance indices $I_{i_{key}}$ and $I_{l_{key}}$ as
$C^{\prime}=\langle\arg\max_{\mathcal{T}_{key}\in\{\mathcal{T}_{i_{key}},\mathcal{T}_{l_{key}}\}}(I_{key}),\mathcal{T}^{1}_{il}\rangle$ (9)
@@
The $C^{\prime}$ denotes the new branch initiated from the key node with the higher importance index. Starting from $\mathcal{T}^{1}_{il}$, the model $\mathcal{L}$ continues conduct the reasoning guided by the information from the new nodes, and since the new nodes introduce fresh perspectives of existing information, the new branch is likely to go and explore previously unconsidered directions with step-by-step reasoning. The process of extending the branch $C^{\prime}$ continues until a specified depth is reached. Eventually, the language model $\mathcal{L}$ recursively generates subsequent nodes, using the strongest connection node as the context for generating new nodes as
$C^{\prime}$ $\rightarrow\mathcal{L}(C^{\prime})=\langle C^{\prime},\mathcal{T}^{2}_{il},\ldots,\mathcal{T}^{K}_{il}\rangle,\quad$ (10)
[…]
Semantic Relevance Selection
@@
Semantic relevance selection is applicable after both key node selection methods. The connection node is selected by the model $\mathcal{L}$ based on semantic relationships between nodes. Then the new branch $C^{\prime}$ continues to extend recursively until a specified depth is reached, with $\mathcal{L}$ generating subsequent nodes based on the strongest semantic relationship as:
@@
$C^{\prime}\rightarrow\mathcal{L}(C^{\prime})=\left\langle T_{\text{selected}},\mathcal{T}_{il}^{1},\mathcal{T}_{il}^{2},\ldots,\mathcal{T}_{il}^{K}\right\rangle$ (11)
@@
where $\mathcal{T}_{\text{selected}}$ is the key node chosen by $\mathcal{L}$, and $K$ denotes the ending depth of the thought chain.
Collaborative Reasoning
@@
Give the task $\mathcal{Q}$ and its unseen complete solution space $\mathcal{P}$, the model $\mathcal{L}$ needs to generate new thought branches based on the original thought structure $\mathcal{S}$. During this process, each new branch $C^{\prime}$ expands the explored thought space $\mathcal{P_{S}}\in\mathcal{P}$ by mining potential solutions based on the established structure: